OpenStudy (anonymous):

prove tan^2(x/2+pi/4) = (1+sinx)/(1-sinx)

OpenStudy (anonymous):

I'll make it easier on you :) Consider : \[\theta = (2x + \pi/4)\] and you'll get : \[= \tan^2 \theta\] simplify this one and in the end plug in the value of theta :) Give it a one more try ^_^